Find the derivative of


F(x)=14+ln  x /√ x + 5




a. 3x+10 / 2 x (x + 5)



b. x +10 / 2x (x + 5)



c. 10 / 2x ( x + 5)



d. 2x (x + 5)


1
Expert's answer
2022-04-28T05:20:56-0400

"F(x)=14+\frac{ln(x)}{\sqrt{x+5}}"

"F'(x)=\frac{(ln(x)'\sqrt{x+5}-(\sqrt{x+5})'ln(x)}{x+5}+0=\frac{\sqrt{x+5}/x- ln(x)/(2\sqrt{x+5})}{x+5}=\frac{1}{x\sqrt{x+5}}-\frac{ln(x)}{2(x+5)^{3/2}}"